# Positive definiteness and Fell bundles over discrete groups

### v1

@@ 6 Some further developments
6.1
Let ${\mathcal{B}}=(B_{g})_{g\in G}$ be a Fell bundle over a discrete group $G$, and let $T$ be a ${\mathcal{B}}$-bundle map. It would be interesting to find some sufficient condition, other than $T$ being a linear combination of positive definite ${\mathcal{B}}$-bundle maps (cf. Remark 3.15), which ensures that $T$ is full or reduced, in which case one would also like to know whether $\Phi_{T}$ or $M_{T}$ is completely bounded. We mention here one result of this type. To any bounded function $\phi:G\to{\mathbb{C}}$ one may associate the ${\mathcal{B}}$-bundle map $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ given by
@@
for every $g\in G$ and $b\in B_{g}$, as we did in Example 4.5 in the case where $\phi$ is positive definite. One may wonder under which conditions $T^{\phi}$ is full or reduced.
Let $B(G)$ denote the Fourier-Stieltjes algebra of $G$ [36], which consists of the coefficient functions of unitary representations of $G$. If $\phi$ belongs to $B(G)$, then $T^{\phi}$ is both full and reduced (because $\phi$ may be written as a linear combination of positive definite functions on $G$). One may also consider coefficient functions of more general representations of $G$ on Hilbert spaces, as was initiated by De Cannière and Haagerup [21] in the case of group bundles. (Note that this will bring something new only when $G$ is non-amenable.) The following result extends [9, Proposition 4.2] (which is itself an extension of [21, Theorem 2.2]) to the setting of Fell bundles.
@@
Let $v$ be a uniformly bounded representation of $G$ into the invertible bounded linear operators on some Hilbert space ${\mathcal{H}}$, and set $K:=\sup\{\|v(g)\|,g\in G\}<\infty$. Let $\xi_{1},\xi_{2}\in{\mathcal{H}}$ and define $\phi:G\to{\mathbb{C}}$ by
@@
It is not difficult to check that there is an adjointable operator $W$ on the Hilbert $B_{e}$-module ${\mathcal{H}}\otimes\ell^{2}({\mathcal{B}})$ satisfying
@@
It is also clear that $W$ is invertible. Then we have
@@
for every $g\in G$ and $b\in B_{g}$. Indeed, for every $\xi\in{\mathcal{H}},h\in G\text{ and~}c\in B_{h}$, we have
$W(I_{\mathcal{H}}\otimes\lambda^{\mathcal{B}}_{g}(b))(\xi\otimes j^{\mathcal{B}}_{h}(c))=W(\xi\otimes j^{\mathcal{B}}_{gh}(bc))=v(gh)\otimes j^{\mathcal{B}}_{gh}(bc)$
$=(v(g)\otimes\lambda^{\mathcal{B}}_{g}(b))(v(h)\otimes j^{\mathcal{B}}_{h}(c))=(v(g)\otimes\lambda^{\mathcal{B}}_{g}(b))W(\xi\otimes j^{\mathcal{B}}_{h}(c)).$
For $\xi\in{\mathcal{H}}$, let $\theta_{\xi}:\ell^{2}({\mathcal{B}})\to{\mathcal{H}}\otimes\ell^{2}({\mathcal{B}})$ be the adjointable operator given by $\theta_{\xi}f=\xi\otimes f$ for all $f\in\ell^{2}({\mathcal{B}})$, whose adjoint is determined by $\theta_{\xi}^{*}(\xi^{\prime}\otimes f)=\langle\xi,\xi^{\prime}\rangle\,f$ for $\xi^{\prime}\in{\mathcal{H}}$ and $f\in\ell^{2}({\mathcal{B}})$. We may now define $M_{\phi}:{\mathcal{L}}_{B_{e}}(\ell^{2}({\mathcal{B}}))\to{\mathcal{L}}_{B_{e}}(\ell^{2}({\mathcal{B}}))$ by
@@
$M_{\phi}(\lambda^{\mathcal{B}}_{g}(b))f=\theta_{\xi_{1}}^{*}W(I_{\mathcal{H}}\otimes\lambda^{\mathcal{B}}_{g}(b))W^{-1}\theta_{\xi_{2}}f=\theta_{\xi_{1}}^{*}(v(g)\otimes\lambda^{\mathcal{B}}_{g}(b))\theta_{\xi_{2}}f$
@@
6.2
In this subsection, $G$ and $H$ are discrete groups, $\varphi\in{\rm Hom}(G,H)$, and $\phi:G\to{\mathbb{C}}$ is a bounded function. We will say that $\phi$ is a reduced $(G,\varphi,H)$-multiplier (or simply a $\varphi$-multiplier, when $G$ and $H$ are clear from the context) if there exists a bounded linear map $M_{\varphi,\phi}:C^{*}_{r}(G)\to C^{*}_{r}(H)$ such that
@@
for every $g\in G$. One can also define full $(G,\varphi,H)$-multipliers in a similar way, in terms of the existence a suitable bounded linear map $\Phi_{\varphi,\phi}:C^{*}(G)\to C^{*}(H)$. Of course, when $G=H$ and $\varphi={\rm id}_{G}$, we recover the usual notions of reduced and full multipliers of $G$ (see e.g. [7] and references therein); the map $M_{{\rm id}_{G},\phi}$ is then denoted $M_{\phi}$, while $\Phi_{{\rm id}_{G},\phi}$ is denoted by $\Phi_{\phi}$. In this case, it is well-known (and follows from Theorems 3.12 and 3.10) that the function $\phi$ is positive definite if and only if $\phi$ is a reduced (resp. full) multiplier and $M_{\phi}$ (resp. $\Phi_{\phi}$) is completely positive.
Let’s rephrase these notions in terms of the group bundles ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ and =(×${\mathcal{B}}=({\mathbb{C}}\times\{h\})_{h\in H}$. Let $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ be the ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map given by
@@
Then it follows readily that $\phi$ is a reduced (resp. full) $(G,\varphi,H)$-multiplier if and only if the bundle map $T^{\phi}$ is reduced (resp. full). Moreover, under the natural identifications $C_{r}^{*}({\mathcal{A}})\simeq C_{r}^{*}(G)$ and $C_{r}^{*}({\mathcal{B}})\simeq C_{r}^{*}(H)$ (resp. $C^{*}({\mathcal{A}})\simeq C^{*}(G)$ and $C^{*}({\mathcal{B}})\simeq C^{*}(H)$), the map $M_{T^{\phi}}$ corresponds to $M_{\phi}$ (resp. $\Phi_{T^{\phi}}$ corresponds to $\Phi_{\phi}$) whenever $\phi$ is a reduced (resp. full) $(G,\varphi,H)$-multiplier.
Note that $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ is positive definite if and only if for all $g_{1},\ldots,g_{n}\in G$ and $z_{1},w_{1},\ldots,z_{n},w_{n}\in{\mathbb{C}}$, we have
$\sum_{i,j=1}^{n}(w_{i},\varphi(g_{i}))\,T^{\phi}_{g_{i}^{-1}g_{j}}\big{(}(z_{i},g_{i})^{*}(z_{j},g_{j})\big{)}\,(w_{j},\varphi(g_{j}))^{*}\,\geq 0\text{ in }{\mathbb{C}}\times\{e\}$
that is,
$\sum_{i,j=1}^{n}w_{i}\bar{w}_{j}\phi(g_{i}^{-1}g_{j})\bar{z}_{i}z_{j}\geq 0\,.$
@@
This is clearly equivalent to $\phi$ being a positive definite function on $G$. Hence, Theorem 3.10 gives that $\phi$ is positive definite if and only $\phi$ is a full $(G,\varphi,H)$-multiplier such that $\Phi_{\varphi,\phi}:C^{*}(G)\to C^{*}(H)$ is completely positive. Moreover, if ${\rm ker}(\varphi)$ is amenable, then Theorem 3.12, gives that this is also equivalent to $\phi$ being a reduced $(G,\varphi,H)$-multiplier such that $M_{\varphi,\phi}:C_{r}^{*}(G)\to C_{r}^{*}(H)$ is completely positive.
@@
The natural problem of determining the space of all reduced (resp. full) $(G,\varphi,H)$-multipliers, and eventually obtain some structure theorem, seems to be challenging. We include below a few relevant observations.
We first note that $\phi$ is a full multiplier of $G$ (i.e., $\phi$ belongs to $B(G)$) if and only if $\phi$ is a full $(G,\theta,K)$-multiplier for every group $K$ and every $\theta\in{\rm Hom}(G,K)$, in which case we have
$\Phi_{\theta,\phi}=\theta_{*}\circ\Phi_{\phi},$
where $\theta_{*}:C^{*}(G)\to C^{*}(K)$ is the canonical *-homomorphism, cf. [11, Corollary 8.C.13]. It would be interesting to know if there are situations where some full $(G,\varphi,H)$-multiplier is not a full multiplier of $G$.
Similarly, $\phi$ is a reduced multiplier of $G$ if and only if it is a reduced $(G,\theta,K)$-multiplier for every group $K$ and every $\theta\in{\rm Hom}(G,K)$ such that $\ker(\theta)$ is amenable, in which case we have
$M_{\theta,\phi}=\theta_{*,r}\circ M_{\phi},$ (7)
where $\theta_{*,r}:C_{r}^{*}(G)\to C_{r}^{*}(K)$ is the canonical *-homomorphism [11, Corollary 8.C.15].
If $\ker(\varphi)$ is not amenable, then it can happen that $\phi$ is a reduced multiplier of $G$, but not a reduced $(G,\varphi,H)$-multiplier. For example, assume $G$ is a non-amenable group, $H=\{e\}$ and $\varphi_{0}:G\to\{e\}$ is the trivial map. Then $\ker(\varphi_{0})$ is not amenable. Moreover, the constant function $1_{G}$ is a reduced multiplier of $G$, but it is not a reduced $(G,\varphi_{0},\{e\})$-multiplier. Indeed, if it was, then $M_{\varphi_{0},1_{G}}$ would correspond to a multiplicative linear functional on $C_{r}^{*}(G)$ taking the value $1$ on each $\lambda_{G}(g)$, which would imply that $G$ is amenable, cf. [14]. It should also be noted here that the space of reduced $(G,\varphi_{0},\{e\})$-multipliers may be identified with the dual of $C_{r}^{*}(G)$.
One may wonder whether there exists some reduced $(G,\varphi,H)$-multiplier which is not a reduced multiplier of $G$, even if $\ker(\varphi)$ is amenable.
Let now $\psi:H\to{\mathbb{C}}$ be a bounded function, so $\psi\circ\varphi:G\to{\mathbb{C}}$ is bounded too. We note that if $\psi$ is a full multiplier of $H$, i.e., $\psi\in B(H)$, then $\psi\circ\varphi$ belongs to $B(G)$, cf. [36, Theorem 2.2.1 (i)]; hence $\psi\circ\varphi$ is a full $(G,\varphi,H)$-multiplier, satisfying that
$\Phi_{\varphi,\psi\circ\varphi}=\Phi_{\psi}\circ\varphi_{*}=\varphi_{*}\circ\Phi_{\psi\circ\varphi}.$
We don’t know if it can happen that $\psi\circ\varphi$ is a full $(G,\varphi,H)$-multiplier for some $\psi\not\in B(H)$.
Further, assume that $\ker(\varphi)$ is amenable and let $\psi$ be a reduced multiplier of $H$. Then one readily sees that $\psi\circ\varphi$ is a reduced $(G,\varphi,H)$-multiplier with
$M_{\varphi,\psi\circ\varphi}=M_{\psi}\circ\varphi_{*,r}.$ (8)
It is not obvious to us that $\psi\circ\varphi$ is necessarily a reduced multiplier of $G$. However, if $M_{\psi}$ is completely bounded, i.e., $\psi\in M_{0}A(H)$, then $\psi\circ\varphi\in M_{0}A(G)$ (this may be deduced from [35, Theorem], which is a variant of Gilbert’s theorem), in which case we get that
$M_{\varphi,\psi\circ\varphi}=M_{\psi}\circ\varphi_{*,r}=\varphi_{*,r}\circ M_{\psi\circ\varphi}.$
Also here, one may wonder whether it can happen that $\psi\circ\varphi$ is a reduced $(G,\varphi,H)$-multiplier for some $\psi$ which is not a reduced multiplier of $H$.
We finally note that if $\psi$ is assumed to be positive definite, then it follows readily that the maps $\Phi_{\varphi,\psi\circ\varphi}$ and $M_{\varphi,\psi\circ\varphi}$ are both completely positive.
6.3
@@
There has been recently a lot of interest in $C^{*}$-algebras associated to Fell bundles, not only over discrete groups but also over locally compact groups, étale groupoids and inverse semigroups (see e.g. [15, 1, 17, 41, 2, 42, 43, 39, 44, 38, 20, 61]). It seems natural to investigate to which extent our results in this paper can be carried over to these settings. Although some work will be necessary in order to handle some technicalities, we expect that most of our results will continue to hold for Fell bundles over locally compact groups. In the other cases, it is likely that some efforts will be needed to formulate the correct versions. For instance, for étale groupoids, one would first have to decide which class(es) of functors between étale groupoids have the property that a suitable generalization of [11, Theorems 8.C.12 and 8.C.14] is achievable. We refer to [61] for some possible candidates.
@@
In another direction, it has been known for quite some time that Fell bundles over (discrete) groupoids and $C^{*}$-categories are close relatives, although it is hard to find any specific statement in the literature. Let ${\cal T}$ be a (small) $C^{*}$-category [30]. One can naturally associate to ${\cal T}$ a Fell bundle ${\mathcal{B}}_{\cal T}$ over the (discrete) pair groupoid ${\rm Ob}({\cal T})\times{\rm Ob}({\cal T})$ by setting $B_{\rho,\sigma}:=(\rho,\sigma)$, the Banach space of arrows from the object $\rho$ to the object $\sigma$ in ${\cal T}$. The bundle operations follow at once from the corresponding operations in ${\cal T}$. If $t\in B_{\rho,\sigma}$ and $s\in B_{\sigma,\tau}$ then $ts\in B_{\rho,\tau}$ (product in the bundle) is given by $s\circ t$ (composition in ${\cal T}$), and $t^{*}\in B_{\sigma,\rho}$ (adjoint in the bundle) is nothing but $t^{*}\in(\sigma,\rho)$. Moreover, it is well known that to any $C^{*}$-category ${\cal T}$ one can associate a $C^{*}$-algebra $C^{*}({\cal T})$ (cf. [30]), and to any Fell bundle ${\mathcal{B}}$ over a groupoid one can associate $C^{*}({\mathcal{B}})$ and $C_{r}^{*}({\mathcal{B}})$ (see e.g. [37]). One may wonder whether there is some relationship between $C^{*}({\cal T})$ and $C^{*}({\mathcal{B}}_{\cal T})$ (or $C_{r}^{*}({\mathcal{B}}_{\cal T})$).

### v2

@@ 6 Some further developments
6.1 Completely bounded maps on cross-sectional C*-algebras
Let ${\mathcal{B}}=(B_{g})_{g\in G}$ be a Fell bundle over a discrete group $G$, and let $T$ be a ${\mathcal{B}}$-bundle map. It seems interesting to look for sufficient conditions ensuring that $T$ is full or reduced with $\Phi_{T}$ or $M_{T}$ completely bounded, cf. Remark 3.17 for a result of this type. Here we present another one.
Given $\phi\in\ell^{\infty}(G)$, let $T^{\phi}$ be the ${\mathcal{B}}$-bundle map given by $T^{\phi}=(T^{\phi}_{g})_{g\in G}$ given by
@@ 6.1 Completely bounded maps on cross-sectional C*-algebras
for every $g\in G$ and $b\in B_{g}$ (as in Example 4.4 when $\phi$ is positive definite).
Let $B(G)$ denote the Fourier-Stieltjes algebra of $G$ [37], consisting of the coefficient functions of unitary representations of $G$. If $\phi$ belongs to $B(G)$, then $T^{\phi}$ is both full and reduced (as $\phi$ is a linear combination of positive definite functions on $G$). One may also consider coefficient functions of more general representations of $G$ on Hilbert spaces, as initiated by De Cannière and Haagerup [21] in the case of group bundles. (Note that this will bring something new only when $G$ is non-amenable.) The following result extends [9, Proposition 4.2] (which is itself an extension of [21, Theorem 2.2]) to the setting of Fell bundles.
@@
Let $v$ be a uniformly bounded representation of $G$ into the invertible bounded linear operators on some Hilbert space ${\mathcal{H}}$, and set $K:=\sup\{\|v(g)\|,g\in G\}<\infty$. Let $\xi_{1},\xi_{2}\in{\mathcal{H}}$ and define $\phi\in\ell^{\infty}(G)$ by
@@
It is not difficult to check that there is an invertible adjointable operator $W$ on the Hilbert $B_{e}$-module ${\mathcal{H}}\otimes\ell^{2}({\mathcal{B}})$ satisfying
@@
A straightforward computation gives that
@@
for every $g\in G$ and $b\in B_{g}$. Now, for $\xi\in{\mathcal{H}}$, let $\theta_{\xi}:\ell^{2}({\mathcal{B}})\to{\mathcal{H}}\otimes\ell^{2}({\mathcal{B}})$ be the adjointable operator given by $\theta_{\xi}f=\xi\otimes f$ for all $f\in\ell^{2}({\mathcal{B}})$, whose adjoint is determined by $\theta_{\xi}^{*}(\xi^{\prime}\otimes f)=\langle\xi,\xi^{\prime}\rangle\,f$ for $\xi^{\prime}\in{\mathcal{H}}$ and $f\in\ell^{2}({\mathcal{B}})$. We may then define $M_{\phi}:{\mathcal{L}}_{B_{e}}(\ell^{2}({\mathcal{B}}))\to{\mathcal{L}}_{B_{e}}(\ell^{2}({\mathcal{B}}))$ by
@@
$~M_{\phi}(\lambda^{\mathcal{B}}_{g}(b))f$ $=\theta_{\xi_{1}}^{*}W(I_{\mathcal{H}}\otimes\lambda^{\mathcal{B}}_{g}(b))W^{-1}\theta_{\xi_{2}}f=\theta_{\xi_{1}}^{*}(v(g)\otimes\lambda^{\mathcal{B}}_{g}(b))\theta_{\xi_{2}}f$
@@
6.2 Group multipliers
Let $G$ and $H$ be discrete groups, $\varphi\in{\rm Hom}(G,H)$, and $\phi:G\to{\mathbb{C}}$ be a bounded function. Let us say that $\phi$ is a reduced $(G,\varphi,H)$-multiplier if there exists a bounded linear map $M_{\varphi,\phi}:C^{*}_{r}(G)\to C^{*}_{r}(H)$ such that
@@ 6.2 Group multipliers
for every $g\in G$. One can also define full $(G,\varphi,H)$-multipliers in a similar way, in terms of the existence a suitable bounded linear map $\Phi_{\varphi,\phi}:C^{*}(G)\to C^{*}(H)$. Of course, when $G=H$ and $\varphi={\rm id}_{G}$, we recover the usual notions of reduced and full multipliers of $G$ (see e.g. [7] and references therein); the map $M_{{\rm id}_{G},\phi}$ is then denoted $M_{\phi}$, while $\Phi_{{\rm id}_{G},\phi}$ is denoted by $\Phi_{\phi}$. In this case, it is well-known (and follows from Theorems 3.14 and 3.11) that the function $\phi$ is positive definite if and only if $\phi$ is a reduced (resp. full) multiplier and $M_{\phi}$ (resp. $\Phi_{\phi}$) is completely positive.
In terms of the group bundles ${\mathcal{A}}=({\mathbb{C}}\times\{g\})_{g\in G}$ and =(×${\mathcal{B}}=({\mathbb{C}}\times\{h\})_{h\in H}$, letting $T^{\varphi,\phi}=(T^{\varphi,\phi}_{g})_{g\in G}$ be the ${\mathcal{A}}$-$\varphi$$-{\mathcal{B}}$ bundle map given by
@@
it is clear that $\phi$ is a reduced (resp. full) $(G,\varphi,H)$-multiplier if and only if the bundle map $T^{\varphi,\phi}$ is reduced (resp. full). Moreover, under the natural identifications $C_{r}^{*}({\mathcal{A}})\simeq C_{r}^{*}(G)$ and $C_{r}^{*}({\mathcal{B}})\simeq C_{r}^{*}(H)$ (resp. $C^{*}({\mathcal{A}})\simeq C^{*}(G)$ and $C^{*}({\mathcal{B}})\simeq C^{*}(H)$), the map $M_{T^{\varphi,\phi}}$ corresponds to $M_{\varphi,\phi}$ (resp. $\Phi_{T^{\varphi,\phi}}$ corresponds to $\Phi_{\varphi,\phi}$) whenever $\phi$ is a reduced (resp. full) $(G,\varphi,H)$-multiplier.
@@
One readily checks that $T^{\varphi,\phi}$ is positive definite if and only if $\phi$ is a positive definite function on $G$. Hence, Theorem 3.11 gives that $\phi$ is positive definite if and only $\phi$ is a full $(G,\varphi,H)$-multiplier such that $\Phi_{\varphi,\phi}:C^{*}(G)\to C^{*}(H)$ is completely positive. Moreover, if ${\rm ker}(\varphi)$ is amenable, then Theorem 3.14, gives that this is also equivalent to $\phi$ being a reduced $(G,\varphi,H)$-multiplier such that $M_{\varphi,\phi}:C_{r}^{*}(G)\to C_{r}^{*}(H)$ is completely positive.
@@
The natural problem of determining the space of all reduced (resp. full) $(G,\varphi,H)$-multipliers seems to be challenging. We hope to return to this point in future work.
6.3 Fell bundles over groupoids and C*-categories
@@
Recently there has been a lot of interest in $C^{*}$-algebras associated to Fell bundles, not only over discrete groups but also over locally compact groups, étale groupoids and inverse semigroups (see e.g. [15, 1, 17, 42, 2, 43, 44, 40, 45, 39, 20, 62]). It seems natural to investigate to which extent our results in this paper can be carried over to these settings. Although some work will be necessary in order to handle some technicalities, we expect that most of our results will continue to hold for Fell bundles over locally compact groups. In the other cases, it is likely that some efforts will be needed to formulate the correct versions. For instance, for étale groupoids, one would first have to decide which class(es) of functors between étale groupoids have the property that a suitable generalization of [11, Theorems 8.C.12 and 8.C.14] is achievable. We refer to [62] for some possible candidates.
@@
In another direction, it has been known for quite some time that Fell bundles over (discrete) groupoids and $C^{*}$-categories are close relatives, although it is hard to find any specific statement in the literature. Let ${\cal T}$ be a (small) $C^{*}$-category [31]. One can naturally associate to ${\cal T}$ a Fell bundle ${\mathcal{B}}_{\cal T}$ over the (discrete) pair groupoid ${\rm Ob}({\cal T})\times{\rm Ob}({\cal T})$ by setting $B_{\rho,\sigma}:=(\rho,\sigma)$, the Banach space of arrows from the object $\rho$ to the object $\sigma$ in ${\cal T}$. The bundle operations follow at once from the corresponding operations in ${\cal T}$. If $t\in B_{\rho,\sigma}$ and $s\in B_{\sigma,\tau}$ then $ts\in B_{\rho,\tau}$ (product in the bundle) is given by $s\circ t$ (composition in ${\cal T}$), and $t^{*}\in B_{\sigma,\rho}$ (adjoint in the bundle) is nothing but $t^{*}\in(\sigma,\rho)$. Moreover, it is well known that to any $C^{*}$-category ${\cal T}$ one can associate a $C^{*}$-algebra $C^{*}({\cal T})$ (cf. [31]), and to any Fell bundle ${\mathcal{B}}$ over a groupoid one can associate $C^{*}({\mathcal{B}})$ and $C_{r}^{*}({\mathcal{B}})$ (see e.g. [38]). A natural guess would be that $C^{*}({\cal T})\simeq C^{*}({\mathcal{B}}_{\cal T})$.
@@
Acknowledgements
We thank the referee for her/his careful reading and for making a long list of helpful comments and suggestions. Erik Bédos is grateful to the Trond Mohn Foundation (TMS) for partial financial support through the project “Pure Mathematics in Norway” during his visit to the Sapienza University of Rome in August 2023. Roberto Conti (RC) visited the University of Oslo (UiO) in June 2023 and in June 2024. Both these stays were partially supported by UiO. RC was also partially supported by Sapienza Università di Roma and INdAM-GNAMPA project Operator algebras and infinite quantum systems CUP E53C23001670001.
7 Statements and Declarations
Competing interests  The authors have no Conflict of interest to declare that are relevant to the content of this article. There are no financial and non-financial conflicts of interest.
Data availability Data sharing is not applicable to this article as no new data have been created or analyzed in this research.